# A Host–Kra $\mathbb{F}_{2}^{\omega}$-system of order $5$ that is not Abramov of order $5$, and non-measurability of the inverse theorem for the $U^{6}(\mathbb{F}_{2}^{n})$ norm

### v2

@@ 5.2. Equidistribution theory for $Q,S$
Observe that this random tuple is nothing more than the image under $\pi_{s}$ of the randomly sampled tuple $(\tilde{Q},\tilde{S})_{(Q,S),x,h_{1},\dots,h_{s}}$. The claim follows. ∎
@@ 5.3. Concluding the argument
The first step is to use Corollary 5.6 to transfer the various structural conclusions of Proposition 5.3 to the nilspace $X_{(Q_{0},S_{0})}$, in order to obtain a situation somewhat resembling that in the proof of Proposition 4.5. Let $n$ be in the indicated sequence going to infinity, and let $x,h_{1},\dots,h_{6}$ be drawn uniformly and independently at random from $\mathbb{F}_{2}^{n}$. By Corollary 5.6, the random tuple
(44) $((a\mapsto(Q,S)((a,0^{n-M})+x+\sum_{i=1}^{6}\omega_{i}h_{i})))_{\omega\in\{0,1\}^{6}}$
(which is a $6$-cube in $X_{(Q_{0},S_{0})}$) converges in distribution to the uniform distribution on $C^{6}(X_{(Q_{0},S_{0})})$, while the random function
@@
on $\mathbb{F}_{2}^{M}$ (which is an element of $X_{(Q_{0},S_{0})}$) converges in distribution to the uniform distribution on $X_{(Q_{0},S_{0})}$.
From (39), (31) we have (for $\varepsilon(m)$ small enough) that
$\mathbb{E}_{x\in\mathbb{F}_{2}^{n}}|e(S(x))-e(\Phi((S(x+a))_{a\in\mathbb{F}_{2}^{M}}))|\geq\frac{1}{2m}.$
Since the random variable (45) converges to the uniform distribution on $X_{(Q_{0},S_{0})}$, we conclude that
(46) $|\mathbb{E}_{(Q^{\prime},S^{\prime})\in X_{(Q_{0},S_{0})}}e(S^{\prime}(0)-\Phi(S^{\prime}))|\geq\frac{1}{2m}.$
@@
Let us assume this proposition for the moment and see how it implies Lemma 5.7. Applying this proposition with $d=0$, we see that for each $q\in X_{2}$, we can find $(Q^{*}_{q},S^{*}_{q})\in X_{(Q_{0},S_{0})}$ obeying the $d=0$ conclusions (1)-(3) of the proposition; in particular, $Q^{*}_{q}(0)=q$. Now let $(q_{\omega})_{\omega\in\{0,1\}^{6}}\in C^{6}(X_{2})$ be a $6$-cube in $X_{2}$, and let $((Q^{\prime}_{\omega},S^{\prime}_{\omega}))_{\omega\in\{0,1\}^{6}}\in C^{6}(X_{(Q_{0},S_{0})})$ be as in the proposition. For each $\omega\in\{0,1\}^{6}$, the point $(Q^{\prime}_{\omega},S^{\prime}_{\omega})$ in $X_{(Q_{0},S_{0})}$ obeys the $d=0$ axioms of (1)-(3) with respect to the 00-cube $q_{\omega}$. Since $(Q^{*}_{q_{\omega}},S^{*}_{q_{\omega}})$ does also, we conclude from the uniqueness component of this proposition that $(Q^{\prime}_{\omega},S^{\prime}_{\omega})\sim(Q^{*}_{q_{\omega}},S^{*}_{q_{\omega}})$ for all $\omega\in\{0,1\}^{d}$. Lemma 5.7 follows.
@@
whenever $11\leq l\leq k-$ 1 and $1\leq i_{1}<\dots\leq i_{l}\leq M$. We claim that the same statement also holds for $l=k$. Indeed, by construction of $X_{(Q_{0},S_{0})}$, we can find $(\tilde{Q},\tilde{S})\in\Sigma^{(1+d)}_{(Q,S)}$ such that
@@
As $(\tilde{Q},\tilde{S})$ is an $M+1+d$-cube in $X_{5,5}$, the right-hand side is equal to
@@
From axiom (3), $02(-S^{\prime}_{\omega}(0)+S^{\prime\prime}_{\omega}(0))=-2S(0)+2S(0)=$ 0, hence for each $\omega\in\{0,1\}^{d}$, $S^{\prime\prime}_{\omega}$ is either equal to $S^{\prime}_{\omega}$ or $S^{\prime}_{\omega}+\frac{1}{2}$. Since also $Q^{\prime}_{\omega}=Q^{\prime\prime}_{\omega}$, we conclude that $(Q^{\prime}_{\omega},S^{\prime}_{\omega})\sim(Q^{\prime\prime}_{\omega},S^{\prime\prime}_{\omega})$. This completes the proof of uniqueness.
Now we establish existence. Let $d\geq 0$, and let $(q_{\omega})_{\omega\in\{0,1\}^{d}}$ be a $d$-cube in $X_{2}$. By the construction of $X_{(Q_{0},S_{0})}$, our task is to find a $M+1+d$-cube $(\tilde{Q},\tilde{S})$ in $X_{5,5}$ obeying the following properties:
@@
    For $a\in\mathbb{F}_{2}^{M}$, we have $\tilde{S}(a,0^{n-M})=S_{0}(a)$. It is not necessarily the case that $\tilde{S}$ agrees with $S$ on $\mathbb{F}_{2}^{M}$ (so that $(\tilde{Q},\tilde{S})$ lies in $\Sigma^{(1+d)}_{(Q,S)}$).
@@
    For every $11\leq l\leq k-$ 1 and $1\leq i_{1}<\dots<i_{l}\leq M$, (49) holds.
@@
First, we use the fact that $\rho=d^{5}\psi$ to write the condition (1) on the $M+1+d$-cube $(\tilde{Q},\tilde{S})$ as
@@
By combining these constructions with the arguments in Appendix B, we obtain a counterexample to Conjecture 1.2. It is natural to ask whether there is a shortcut approach that could construct the counterexample to Conjecture 1.2 more directly, without first building a counterexample to Conjecture 1.3. Morally speaking, this should proceed by starting with the space $\operatorname{Hom}_{\Box}({\mathcal{D}}^{1}(\mathbb{F}_{2}^{\omega})\to X_{5,5})$, which is a compact $\mathbb{F}_{2}^{\omega}$-system that can be naturally equipped with a Haar measure. This system is not ergodic, but a generic component of the ergodic decomposition should be a $5$-step ergodic $\mathbb{F}_{2}^{\omega}$-system that fails to be Abramov of order $5$ (cf., the role of the pair $(Q_{0},S_{0})$ in the above analysis); However, in the spirit of Remark 4.6, Candela et al [5] recently established that this system admits an extension that is Abramov of order $5$ (and it should even be a Weyl system of order $5$, in the sense of e.g., [18]). The rigorous verification of these claims seems to be of comparable complexity to the arguments just presented, and so we do not detail this more direct approach here.
@@ Appendix A Nilspaces, filtered abelian groups, and non-classical polynomials
In this section, we gather several standard (and primarily algebraic) facts about nilspaces, filtered abelian groups, and polynomial maps. Most of the concepts introduced below are not new. However, different authors often use different terminology for the same concept, which can lead to confusion. For example, the notion of a $k$-cocycle (see Definition A.6) was originally introduced by Gutman, Manners, and Varjú (see [15, Definition 4.8]). The same concept was also defined by Candela in [3], where it is referred to as a "degree $k$-cocycle." To avoid confusion between the degree of a polynomial and that of a cocycle, we have chosen to adopt the terminology of Gutman, Manners, and Varjú. Similarly, what we refer to in this paper as $k$-coboundaries, and $k$-extensions are called degree $k$ coboundaries, and degree $k$ extensions by Candela in [3]. It is also important to emphasize that our concept of a $k$-cocycle is not the same as the similarly named concept in homological algebra, where $k$-cocycles are $k$-cochains with vanishing coboundary (in fact, from the perspective of homological algebra all our cocycles are $1$-cocycles).
@@ A.1. Nilspaces
Nilspaces were introduced by Host–Kra [17] (under the equivalent formulation of parallelepiped structures) and Camarena–Szegedy [2] as an abstraction of the concept of a parallelepiped in a group or dynamical system. They can be defined in the set-theoretic, topological, and measurable categories, but we will only need to consider finite nilspaces, which allows us to work in the technically simpler set-theoretic category. We recall the definition of a nilspace, following [3, Definition 1.2.1]:
@@
    (00-ergodicity) $C^{0}(X)=X$. If we have the stronger property $C^{1}(X)=X^{\{0,1\}}$, we say that the nilspace is ergodic (or $1$-ergodic).
@@
    (Corner completion) Let $n\geq 1$, and let $c^{\prime}\colon\{0,1\}^{n}\backslash\{1\}^{n}\to X$ be a function such that every restriction of $c^{\prime}$ to an $(n-1)$-face containing $0^{n}$ is in $C^{n-1}(X)$. Then there exists $c\in C^{n}(X)$ such that $c(v)=c^{\prime}(v)$ for all $v\neq 1^{n}$. If this $c$ is unique, we say that $X$ is an $(n-1)$-step nilspace.
@@
In much of the literature (e.g., [3]) the term “nilspace” is used to denote what we call an “ergodic nilspace”, but it will be convenient for us to only impose the weaker axiom of 00-ergodicity in our basic definitions. In any event, it is often not difficult to reduce to the ergodic case via the following ergodic decomposition. If $X$ is a nilspace, we can define a relation [^5] This is a special case of a more general class of equivalence relations $\sim_{k}$ one can define on nilspaces; see [3, Definition 3.2.3]. $\sim_{0}$ on $X$ by declaring $x\sim_{0}y$ if $(x,y)\in C^{1}(X)$. It is not difficult to verify that this is an equivalence relation, that each equivalence class has the structure of an ergodic nilspace, and the original nilspace $X$ is the disjoint union of these ergodic nilspaces; see [3, Lemma 3.1.8]. Because of this, many of the foundational results on ergodic nilspaces (such as those set out in [3]) extend without difficulty to the more general nilspace setting.
@@
[15, Definition 4.8] Let $X$ be a nilspace, $Z$ be an abelian group, and $k\geq 0$. A $k$-cocycle on $X$ taking values in $Z$ is a function $\rho\colon C^{k+1}(X)\to Z$ obeying the following axioms:
@@
    (Cocycle) If $x,y,z\in C^{k}(X)$ are $k$-cubes with $(x,y),(y,z)\in C^{1}(C^{k}(X))\equiv C^{k+1}(X)$ are $k+1$-cubes (which implies that $(x,z)$ is also a $k+1$-cube, thanks to Remark A.3), then
@@ A $1$-cocycle is a map $\rho\colon C^{2}(X)\to Z$ obeying the symmetry axiom
Now we introduce a key construction.
@@
If $Z$ is an abelian group, then the $k$-step nilspace ${\mathcal{D}}^{k}(Z)$ (defined in the next section) can be thought of as the skew product $\mathrm{pt}\rtimes^{(k)}_{0}Z$ of a point $\mathrm{pt}$ and $Z$ by the zero cocycle 00.
@@
If $\rho=d^{k+1}F$ is a $k$-coboundary, then the skew product $X\rtimes^{(k)}_{\rho}Z$ is isomorphic as a nilspace to the product nilspace $X\times{\mathcal{D}}^{k}(Z)=X\rtimes_{0}Z$, with the isomorphism defined by mapping $(x,z)$ to $(x,z-F(x))$. More generally, adding or subtracting a $k$-coboundary from a cocycle does not affect the skew product up to nilspace isomorphism.
@@ A.2. Filtered abelian groups
(see e.g., [13, §6]) A filtered abelian group $G=(G,(G_{i})_{i\geq 0})$ is an abelian group $G=(G,+)$ (which we will usually think of as being discrete), equipped with a filtration
@@
See [14, Theorem B.3, Proposition B.8] or [3, Theorem 2.2.14] (the latter statement is written in the ergodic case, but the proof extends without difficulty to the non-ergodic setting). ∎
As one corollary of this lemma, we see that the space $\operatorname{Hom}_{\Box}(G\to H)$ of nilspace morphisms from one filtered abelian group $G$ to another $H$ is an abelian group, which contains the space of filtered homomorphisms from $G$ to $H$ as a subgroup. In fact $\operatorname{Hom}_{\Box}(G\to H)$ naturally has the structure of a filtered abelian group, in a manner consistent with the nilspace structure on $\operatorname{Hom}_{\Box}(G\to H)$ already constructed in Remark A.3: see [14, Proposition B.6]. The translation operators $x\mapsto x+h$ on $G$ are also nilspace morphisms for any $h\in H$.

### v3

@@ 5.2. Equidistribution theory for $Q,S$
The distribution of the random tuple above coincides with the distribution $\mu^{(s)}_{Q,S}$ of the random cube $(\tilde{Q},\tilde{S})_{(Q,S),x,h_{1},\dots,h_{s}}$. By definition of $\pi_{s}$, this tuple is exactly the image of $\mu^{(s)}_{Q,S}$ under $\pi_{s}$. Similarly, the uniform distribution on $C^{s}(X_{(Q_{0},S_{0})})$ is the image of $\overline{\mu}^{(s)}_{Q_{0},S_{0}}$ under $\pi_{s}$. Since Corollary 5.5 asserts that $\mu^{(s)}_{Q,S}$ converges in total variation to $\overline{\mu}^{(s)}_{Q_{0},S_{0}}$, applying $\pi_{s}$ yields the desired convergence in distribution of the tuple to uniform on $C^{s}(X_{(Q_{0},S_{0})})$. ∎
@@ 5.3. Concluding the argument
The first step is to use Corollary 5.6 to transfer the structural conclusions of Proposition 5.3 to the nilspace $X_{(Q_{0},S_{0})}$, in a form resembling the proof of Proposition 4.5. Let $n$ belong to the sequence from Corollary 5.5, and let $x,h_{1},\dots,h_{6}$ be chosen uniformly and independently at random from $\mathbb{F}_{2}^{n}$. By Corollary 5.6, the random $6$-cube
(47) $\bigl{(}(a\mapsto(Q,S)((a,0^{n-M})+x+\sum_{i=1}^{6}\omega_{i}h_{i}))\bigr{)}_{\omega\in\{0,1\}^{6}}$
converges in distribution to the uniform distribution on $C^{6}(X_{(Q_{0},S_{0})})$, while the random element
@@
converges in distribution to the uniform distribution on $X_{(Q_{0},S_{0})}$.
On the other hand, combining (42) and (34) gives (for $\varepsilon(m)$ sufficiently small)
$\mathbb{E}_{x\in\mathbb{F}_{2}^{n}}\bigl{|}e(S(x))-e\bigl{(}\Phi((S(x+a))_{a\in\mathbb{F}_{2}^{M}})\bigr{)}\bigr{|}\;\geq\;\frac{1}{2m}.$
[…]
The function $(Q^{\prime},S^{\prime})\mapsto|e(S^{\prime}(0))-e(\Phi(S^{\prime}))|$ is bounded and continuous on the compact nilspace $X_{(Q_{0},S_{0})}$. Since (48) converges in distribution to the uniform distribution on $X_{(Q_{0},S_{0})}$ (by Corollary 5.6), the Portmanteau theorem implies that this expectation converges to its uniform counterpart. Hence
(49) $\Bigl{|}\mathbb{E}_{(Q^{\prime},S^{\prime})\in X_{(Q_{0},S_{0})}}e(S^{\prime}(0)-\Phi(S^{\prime}))\Bigr{|}\;\geq\;\frac{1}{2m}.$
@@
Let us assume this proposition for the moment and see how it implies Lemma 5.7. Applying this proposition with $d=0$, we see that for each $q\in X_{2}$, we can find $(Q^{*}_{q},S^{*}_{q})\in X_{(Q_{0},S_{0})}$ obeying the $d=0$ conclusions (1)-(3) of the proposition; in particular, $Q^{*}_{q}(0)=q$. Now let $(q_{\omega})_{\omega\in\{0,1\}^{6}}\in C^{6}(X_{2})$ be a $6$-cube in $X_{2}$, and let $((Q^{\prime}_{\omega},S^{\prime}_{\omega}))_{\omega\in\{0,1\}^{6}}\in C^{6}(X_{(Q_{0},S_{0})})$ be as in the proposition. For each $\omega\in\{0,1\}^{6}$, the point $(Q^{\prime}_{\omega},S^{\prime}_{\omega})$ in $X_{(Q_{0},S_{0})}$ obeys the $d=0$ axioms of (1)-(3) with respect to the 0-cube $q_{\omega}$. Since $(Q^{*}_{q_{\omega}},S^{*}_{q_{\omega}})$ does also, we conclude from the uniqueness component of this proposition that $(Q^{\prime}_{\omega},S^{\prime}_{\omega})\sim(Q^{*}_{q_{\omega}},S^{*}_{q_{\omega}})$ for all $\omega\in\{0,1\}^{d}$. Lemma 5.7 follows.
@@
whenever $1\leq l\leq k-1$ and $1\leq i_{1}<\dots\leq i_{l}\leq M$. We claim that the same statement also holds for $l=k$. Indeed, by construction of $X_{(Q_{0},S_{0})}$, we can find $(\tilde{Q},\tilde{S})\in\Sigma^{(1+d)}_{(Q,S)}$ such that
@@
[…]
@@
From axiom (3), $2(-S^{\prime}_{\omega}(0)+S^{\prime\prime}_{\omega}(0))=-2S(0)+2S(0)=0$, hence for each $\omega\in\{0,1\}^{d}$, $S^{\prime\prime}_{\omega}$ is either equal to $S^{\prime}_{\omega}$ or $S^{\prime}_{\omega}+\frac{1}{2}$. Since also $Q^{\prime}_{\omega}=Q^{\prime\prime}_{\omega}$, we conclude that $(Q^{\prime}_{\omega},S^{\prime}_{\omega})\sim(Q^{\prime\prime}_{\omega},S^{\prime\prime}_{\omega})$. This completes the proof of uniqueness.
[…]
@@
[…]
@@
    For every $1\leq l\leq k-1$ and $1\leq i_{1}<\dots<i_{l}\leq M$, (52) holds.
@@
[…]
@@
By combining these constructions with the arguments in Appendix B, we obtain a counterexample to Conjecture 1.2. It is natural to ask whether there is a shortcut approach that could construct the counterexample to Conjecture 1.2 more directly, without first building a counterexample to Conjecture 1.3. Morally speaking, this should proceed by starting with the space $\operatorname{Hom}_{\Box}({\mathcal{D}}^{1}(\mathbb{F}_{2}^{\omega})\to X_{5,5})$, which is a compact $\mathbb{F}_{2}^{\omega}$-system that can be naturally equipped with a Haar measure. This system is not ergodic, but a generic component of the ergodic decomposition should be a $5$-step ergodic $\mathbb{F}_{2}^{\omega}$-system that fails to be Abramov of order $5$ (cf., the role of the pair $(Q_{0},S_{0})$ in the above analysis). The rigorous verification of these claims seems to be of comparable complexity to the arguments just presented, and so we do not detail this more direct approach here.
On the other hand, in the spirit of Remark 4.6, Candela et al. [4] recently established that this system corresponding to the generic component of the ergodic decomposition admits an extension that is Abramov of order 55 (and it should even be a Weyl system of order $5$, in the sense of, e.g., [19]).
@@ Appendix A Nilspaces, filtered abelian groups, and non-classical polynomials
In this section, we gather several standard (and primarily algebraic) facts about nilspaces, filtered abelian groups, and polynomial maps. Most of the concepts introduced below are not new, but the terminology varies between authors. For example, the notion of a $k$-cocycle (see Definition A.6) was first introduced by Antolín-Camarena and Szegedy [2, Definition 2.14] under the name of “degree $k$ cocycle”. The same concept was subsequently treated by Candela [3], and later by Gutman, Manners, and Varjú [16, Definition 4.8], who use the terminology “$k$-cocycle”. To avoid confusion between the degree of a polynomial and that of a cocycle, we follow the terminology of Gutman–Manners–Varjú in this paper. We also note that our Definition A.6 is essentially the same as that in [2], whereas the version in [16] includes an additional continuity requirement.
Similarly, the notions of $k$-coboundaries and $k$-extensions, as we use them here, were originally called degree $k$ coboundaries and degree $k$ extensions in [2], and were further developed in Candela [3].
Finally, we emphasize that our concept of a $k$-cocycle is not the same as the similarly named notion from homological algebra, where $k$-cocycles are $k$-cochains with vanishing coboundary (in fact, from the homological perspective all our cocycles are $1$-cocycles).
@@ A.1. Nilspaces
Host and Kra [18] introduced a combinatorial framework for cubes in abstract sets in dimensions $2$ and $3$ as an abstraction of the concept of a parallelepiped in a group or dynamical system, and Antolín-Camarena and Szegedy [2] later extended this to all dimensions, thereby defining the general notion of abstract nilspaces. They can be defined in the set-theoretic, topological, and measurable categories, but we will only need to consider finite nilspaces, which allows us to work in the technically simpler set-theoretic category. We recall the definition of a nilspace, following [3, Definition 1.2.1]:
@@
    (0-ergodicity) $C^{0}(X)=X$. If we have the stronger property $C^{1}(X)=X^{\{0,1\}}$, we say that the nilspace is ergodic (or $1$-ergodic).
@@
[…]
@@
In much of the literature (e.g., [3]) the term “nilspace” is used to denote what we call an “ergodic nilspace”, but it will be convenient for us to only impose the weaker axiom of 0-ergodicity in our basic definitions. In any event, it is often not difficult to reduce to the ergodic case via the following ergodic decomposition. If $X$ is a nilspace, we can define a relation [^11] This is a special case of a more general class of equivalence relations $\sim_{k}$ one can define on nilspaces; see [3, Definition 3.2.3]. $\sim_{0}$ on $X$ by declaring $x\sim_{0}y$ if $(x,y)\in C^{1}(X)$. It is not difficult to verify that this is an equivalence relation, that each equivalence class has the structure of an ergodic nilspace, and the original nilspace $X$ is the disjoint union of these ergodic nilspaces; see [3, Lemma 3.1.8]. Because of this, many of the foundational results on ergodic nilspaces (such as those set out in [3]) extend without difficulty to the more general nilspace setting.
@@
[2, Definition 2.14] Let $X$ be a nilspace, $Z$ be an abelian group, and $k\geq 0$. A $k$-cocycle on $X$ taking values in $Z$ is a function $\rho\colon C^{k+1}(X)\to Z$ obeying the following axioms:
@@
    (Cocycle) If $x,y,z\in C^{k}(X)$ are $k$-cubes with [^13] We denote by $(x,y)$ the $k+1$-cube which is the concatenation of the $k$-cubes $x,y$. $(x,y),(y,z)\in C^{1}(C^{k}(X))\equiv C^{k+1}(X)$ are $k+1$-cubes (which implies that $(x,z)$ is also a $k+1$-cube, thanks to Remark A.3), then
@@ A $1$-cocycle is a map $\rho\colon C^{2}(X)\to Z$ obeying the symmetry axiom
Now we introduce a key construction which originated from [2, Proposition 3.1 and (5)].
@@
If $Z$ is an abelian group, then the $k$-step nilspace ${\mathcal{D}}^{k}(Z)$ (defined in the next section) can be thought of as the skew product $\mathrm{pt}\rtimes^{(k)}_{0}Z$ of a point $\mathrm{pt}$ and $Z$ by the zero cocycle 0.
@@
If $\rho=d^{k+1}F$ is a $k$-coboundary, then the skew product $X\rtimes^{(k)}_{\rho}Z$ is isomorphic as a nilspace to the product nilspace $X\times{\mathcal{D}}^{k}(Z)=X\rtimes_{0}Z$, with the isomorphism defined by mapping $(x,z)$ to (x,zF(x))(x,z-F(x)). More generally, adding or subtracting a $k$-coboundary from a cocycle does not affect the skew product up to nilspace isomorphism.
@@ A.2. Filtered abelian groups
(see e.g., [14, §6]) A filtered abelian group $G=(G,(G_{i})_{i\geq 0})$ is an abelian group $G=(G,+)$ (which we will usually think of as being discrete), equipped with a filtration
@@
See [15, Theorem B.3, Proposition B.8] or [3, Theorem 2.2.14] (the latter statement is written in the ergodic case, but the proof extends without difficulty to the non-ergodic setting). ∎
As one corollary of this lemma, we see that the space $\operatorname{Hom}_{\Box}(G\to H)$ of nilspace morphisms from one filtered abelian group $G$ to another $H$ is an abelian group, which contains the space of filtered homomorphisms from $G$ to $H$ as a subgroup. In fact $\operatorname{Hom}_{\Box}(G\to H)$ naturally has the structure of a filtered abelian group, in a manner consistent with the nilspace structure on $\operatorname{Hom}_{\Box}(G\to H)$ already constructed in Remark A.3: see [15, Proposition B.6]. The translation operators $x\mapsto x+h$ on $G$ are also nilspace morphisms for any $h\in H$.